OpenStudy (anonymous):

lim as x approaches infinity > 3x/(sqrt x^2-1)

hartnn (hartnn):

can you factor out 'x' from the denominator ?

hartnn (hartnn):

or can you factor, x^2 from x^2-1 ?

OpenStudy (anonymous):

x^2-1 is under radical in the denom. can't remember how to solve this

hartnn (hartnn):

ok, so does this make sense ? \((x^2-1)= x^2 (1-1/x^2)\)

OpenStudy (anonymous):

yes, it is still under radical. right?

hartnn (hartnn):

yup. so now \(\sqrt{x^2 (1-1/x^2)}=....?\) now can you factor out x ?

OpenStudy (anonymous):

i know the answer is 3, but can't get the steps.

OpenStudy (anonymous):

does 1/3^2 = 0

hartnn (hartnn):

\(\sqrt{x^2 (1-1/x^2)}=x\sqrt{1-1/x^2}\) got that ^ ?

OpenStudy (anonymous):

got that

hartnn (hartnn):

cool, so what gets cancelled ?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

plug in 3 now?

hartnn (hartnn):

wait, x -> infinity! so, 1/x -> 0 does this make sense ?

OpenStudy (anonymous):

no. i have 3/(x-(1/x^2))

OpenStudy (anonymous):

i don't know how to use the symbols on my computer for this

hartnn (hartnn):

but those 'x' got cancelled from numerator and denominator, right ? so you r left with \(\dfrac{3}{\sqrt {1-1/x^2}}\) isn't it ?

OpenStudy (anonymous):

yep

hartnn (hartnn):

so, just plug in 1/x^2 = 0 in that! what u get ?

OpenStudy (anonymous):

why do you plug in 0?

hartnn (hartnn):

oh, i just said as x-> infinity so 1/x -> 0 so, 1/x^2 will also ->0 so we put 1/x^2 = 0 to get the limit :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^